OpenStudy (anonymous):

the sales force a bussiness consists 10 men and 10 women. a production unit of 7 people are choosen at ramdon. what is the prob it will include 2 men and 5 women?

OpenStudy (anonymous):

didnt I already answer this? do you know permutations and combinations?

OpenStudy (anonymous):

I meant do you understand permutations and combinations?

OpenStudy (anonymous):

Go back to previous time you asked this, I gave more explanation on dhatraditya's work

OpenStudy (anonymous):

now wat

OpenStudy (anonymous):

This is probably the 10th question on this subject you've asked about, and it doesn't seem like you're any closer to learning what you need to know to solve these on your own.. Maybe you should actually spend some time reading the book? \[\frac{{10 \choose 2} * {10 \choose 5}}{20 \choose 7}\]

OpenStudy (anonymous):

i got that up to the ending but then i come w the decinmal

OpenStudy (anonymous):

yes, probabilities are between 0 and 1 and typically decimals

OpenStudy (anonymous):

(10C2*10C5)/(20C7) = 45*252/77520= 1134/7752 = 567/3876 = 189/1

OpenStudy (anonymous):

189/1? how did you get that?

OpenStudy (anonymous):

and then do u devide all them ?

OpenStudy (anonymous):

i have no idea what you are talking about.

OpenStudy (anonymous):

how do u get the answer?

OpenStudy (anonymous):

What you have above is correct except the last fraction.

OpenStudy (anonymous):

ok now what?

OpenStudy (anonymous):

the end? If you want a decimal approximation you could plug that fraction into your calculator

OpenStudy (anonymous):

all them?

OpenStudy (anonymous):

they're all the same fraction (except the 189/1 which I have no idea where it came from)

OpenStudy (anonymous):

so i would devide all those numbers? or?

OpenStudy (anonymous):

1134/7752. use your calculator to do that division. the answer will be a decimal number. that is your answer.

OpenStudy (anonymous):

6.835978835

OpenStudy (anonymous):

so would the answer be 6.84?

OpenStudy (anonymous):

oh come on! what is 1134 divided by 7752?

OpenStudy (anonymous):

0.1462848